OpenStudy (anonymous):

Find the derivative of the exponential function: e^-x^2

OpenStudy (anonymous):

\[e ^{-x ^{2}}\]

OpenStudy (anonymous):

-2xe^(-x^2)

OpenStudy (anonymous):

\[-2\times x \times \exp ^{-x^2}\]

OpenStudy (anonymous):

The second and third responses are both correct. Here's why: \[y=e ^{-x ^{2}}\] is what you gave; if you set a new variable \[u=(-x)^{2}\] Then it's relatively easy to find du/dx: \[du/dx=2(-x)^{2-1}=2(-x)=-2x\] This will come in handy later. Now, if we take the original equation and differentiate it in terms of x, we get \[dy/dx=d/dx(e ^{-x ^{2}})=d/dx(e ^{u})\] By the Chain Rule, we get \[dy/dx=d/dx(e ^{u})=d/du(e ^{u})du/dx\] By differentiating in terms of u, we get \[dy/dx=e ^{u}du/dx\] If we then substitute u and du/dx for the equations we stated above, we get \[dy/dx=e ^{-x ^{^{2}}}(-2x)=-2xe ^{-x ^{2}}\]